OpenStudy (anonymous):

Help with 2 questions? giving you a medal! A business has $15,000 to spend on airline tickets to travel to a conference. It wants 27 of its employees to attend. The business wants to buy as many business-class seats as possible. The business-class seats cost $700. The economy-class seats cost $375. How many of each type of ticket should the business purchase? 11 business class and 16 economy class 12 business class and 15 economy class 15 business class and 12 economy class 16 business class and 11 economy class Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5

OpenStudy (ranga):

First Problem: Assume the number of business class seats purchased to be x There are 27 employees. So what is the number of economy class seats purchased?

OpenStudy (anonymous):

27 class seats ? :O

OpenStudy (ranga):

For 27 employees they need to buy 27 seats. x of the seats are business class. So the remaining seats has to be the economy class. So you need to subtract x from 27. Number of economy class seats = (27 - x) Number of business class seats = x One business class seat costs $700 How much would x business class seat cost?

OpenStudy (anonymous):

i subtracted 700 from 27 and i got 673

OpenStudy (ranga):

No. If one business class seat costs $700 Then x business class seats will cost 700x One economy class seat costs $375 (27 - x) economy class seats will cost 375(27 - x) Total cost is 700x + 375(27 - x) = 15000 Can you solve the above equation for x?

OpenStudy (anonymous):

I think so, hold on

OpenStudy (anonymous):

x=15?

OpenStudy (ranga):

Good! So number of business class seats = x = 15. Number of economy class seats = ?

OpenStudy (anonymous):

can you write the equation for me to solve it ? the way i solved for x!

OpenStudy (ranga):

In one of my replies above it says Number of economy class seats = (27 - x) You have solved for x. So number of economy class seats = ?

OpenStudy (anonymous):

economy class seats is 12? :o

OpenStudy (ranga):

Yes. Business class = 15 and Economy class = 12.

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

Can you help me with question 2?

OpenStudy (anonymous):

and can you help with this one too? How many solutions will this system of equations have? y = -3.5x - 3.5 y = -3.5x + 3.5

OpenStudy (ranga):

For question 2: Start with the given equation 4x + 2y = 9 and divide each term by 2 Can you tell me what the equation will look like if each term is divided by 2?

OpenStudy (anonymous):

OKay so i divided 4 to each side. 4x crossed out so now i have 2y over 4 equals 9 over 4. Where do i go from there?

OpenStudy (ranga):

No. 4x + 2y = 9 Take each term (that is 4x, 2y, 9) and divide it by 2 and tell me what the equation will look like?

OpenStudy (ranga):

If I divide each term by 2 the equation will be: 4x/2 + 2y/2 = 9/2 Can you simplify the above equation?

OpenStudy (anonymous):

4x/2= 2? 2y/2=1? 9/2=4.5? Am i on the right track?

OpenStudy (ranga):

It is: 2x + y = 4.5 Compare this equation to the 4 answers and tell me which one looks wrong.

OpenStudy (anonymous):

x-2y=7 looks wrong

OpenStudy (ranga):

No. The second answer says: 2x + y = -6 We are getting 2x + y = 4.5 The left hand sides are both the same. But the right hand sides are different. And that cannot happen. How can the same (2x + y) be -6 and also be 4.5? So the second answer causes inconsistency and so the answer is the second choice.

OpenStudy (anonymous):

Ohhhhh! thankyou so much, i understand it noww

OpenStudy (ranga):

The third problem has no solutions: y = -3.5x - 3.5 y = -3.5x + 3.5